OpenStudy (anonymous):

Solve by using substitution. 6x + 6y = -6 5x + y = -13 What is the value of y?

OpenStudy (anonymous):

no

pooja195 (pooja195):

if you subtract 5x so that you can get y=-5x-13 you can plug it in

pooja195 (pooja195):

\[6x + 6(-5x-13) = -6\]

OpenStudy (anonymous):

still dont get it lol can someon go in debt more

OpenStudy (anonymous):

True

pooja195 (pooja195):

The method of solving "by substitution" works by solving one of the equations (you choose which one) for one of the variables (you choose which one), and then plugging this back into the other equation, "substituting" for the chosen variable and solving for the other. Then you back-solve for the first variable.

pooja195 (pooja195):

\[6x+6(−5x−13)=−6 \] use the distributive property to solve this :)

pooja195 (pooja195):

http://www.purplemath.com/modules/systlin4.htm <----if you still dont get it go here

OpenStudy (anonymous):

thx so much @pooja195

pooja195 (pooja195):

you're welcome :) shall we finish solving the problem? ;)

OpenStudy (anonymous):

yes

pooja195 (pooja195):

ok so solve what i gave you using the distributive property

OpenStudy (anonymous):

how

pooja195 (pooja195):

\[6x+6(−5x−13)=−6\] whats \[6\times-5?\] whats\[6\times-13?\]

OpenStudy (anonymous):

6 * -5=-30 and 6*-13=-78

pooja195 (pooja195):

good so we have \[6x-30x-78=-6\] combine like terms

OpenStudy (anonymous):

-24x

pooja195 (pooja195):

yes so\[-24x-78=-6\] add -78 to -6 so \[-6+78\]

OpenStudy (anonymous):

72

pooja195 (pooja195):

now do 72/-24

OpenStudy (anonymous):

-3

pooja195 (pooja195):

right so x=-3 do you know what to do next?

OpenStudy (anonymous):

no lol

pooja195 (pooja195):

ok now we have to plug that in our equation WE made :) to get the Y value we had \[y= −5(-3)−13\] solve

OpenStudy (anonymous):

-15

pooja195 (pooja195):

-15-13=?

OpenStudy (anonymous):

-28

pooja195 (pooja195):

wait we have this wrong

pooja195 (pooja195):

\[−5(-3)−13\]

OpenStudy (anonymous):

i got -15 again

pooja195 (pooja195):

no a - + - = +

OpenStudy (anonymous):

15

pooja195 (pooja195):

15-13

OpenStudy (mtalhahassan2):

well done pooja

OpenStudy (anonymous):

2

OpenStudy (confluxepic):

That's correct.

OpenStudy (anonymous):

thank you

OpenStudy (confluxepic):

You can make sure it's correct by plugging in the values and seeing if they make sense.

OpenStudy (anonymous):

ok

pooja195 (pooja195):

Sorry the page froze on me ;-; \(\huge\color{red}{WELCOME~TO~OPENSTUDY!!}\)

OpenStudy (anonymous):

thx

OpenStudy (confluxepic):

@Cooper.dalen14 You may want to read these. http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions Have a good time openstudying. :)